# Continual Adversarial Defense

### 15 Dec 2023 v1

@@ 4.3 Prototype Augmentation for Memory Efficiency
Input: Benign dataset $\mathcal{D}_{\mathrm{train}}$, initial attack $A_{0}$, target model $f_{t}$, and few-shot feedback $\mathcal{A}_{\mathrm{train},K}^{i}$ generated by attack $A_{i}$ for stage $i=1,\cdots,T$.
@@
4.4 Model Ensemble for Clean and Adversarial Image Classification
The overall algorithm of CAD framework is presented in Algorithm 1.
@@ 5 Experiment
To validate the performance of CAD against various adversarial attack methods in the dynamic scenario, we conduct extensive empirical studies on two datasets and compare CAD to baselines from two research streams. For evaluation and analysis, we use two metrics in this section: (1) classification accuracy against each attack after adaptation. (2) the average of the accuracy of all attacks that have occurred in and before in each stage, i.e. average incremental accuracy [28].
@@
We select 11 adversarial attack methods PGD [26], FGSM [6], BIM [14], RFGSM [35], MIM [3], DIM [40], NIM [17], SNIM [17], VNIM [38], and VMIM [38] to compose the attack pool. The perturbation magnitude of all the attacks is set to $\epsilon=8$ under $l_{\infty}$ following [44].
@@
Both the target model and defense model use ResNet as the backbone network in our method. Specifically, we use WideResNet-28-10 [42] for CIFAR-10 following [7] and ResNet-50 [8] for ImageNet-100 following [15]. A 4-layer ConvNet is used as the weight estimator model. The cosine classifier [5] is adopted as the classifier of the defense model. We utilize Torchattacks [12] for generating adversarial images. The amount of feedback per class is set to $K=10$. The number of stages is set to $T=9$ since there are 10 attacks (PGD is the initial attack). Parameters of Beta distribution are set to $\alpha=\beta=2$ and the trade-off parameter for $\mathcal{L}_{0}$ is set to $\gamma=0.01$ following [48]. Epochs for training $f_{0}$ are set to $100$ and fine-tuning epochs are set to $4$.
Baselines
We first compare to adversarial training methods TRADES [44], JEM [7], GAIRAT [45], FastAdv [11], RPF [2] and DMAT [39]. Besides, we propose another variant of our method denoted as premium-CAD in the case of no constraint of memory and no concerns about the amount of feedback. In premium-CAD, we train additional attack-specific defense models using abundant feedback (i.e., 1000 images per class ) in each stage, and ensemble them with the target model by the weight estimator. Details about premium-CAD are given in the supplementary.
Stage 0 1 2 3 4 5 6 7 8 9 -
Attack PGD VNIM FGSM BIM RFGSM MIM DIM NIM SNIM VMIM Clean
TRADES [44] 55.8 57.4 60.7 55.9 55.8 57.1 57.1 58.9 60.2 57.3 84.8
JEM [7] 55.8 46.5 60.7 52.3 52.4 51.1 51.4 57.0 62.0 48.2 91.7
PAT [15] 45.5 43.8 49.8 45.5 45.3 45.2 45.2 43.9 45.1 45.4 69.3
GAIRAT [45] 66.9 66.4 68.3 66.9 66.7 66.5 66.5 68.6 70.4 66.4 89.4
FastAdv [11] 37.1 36.8 41.9 37.0 37.7 37.7 54.6 36.4 39.6 37.6 75.2
RPF [2] 60.9 52.2 58.0 58.9 54.9 57.6 57.5 53.5 52.8 52.6 83.5
DMAT [39] 71.7 73.1 75.8 71.7 71.7 73.0 73.0 75.5 76.0 73.0 92.4
EBM [9] 75.2 73.0 70.0 76.6 76.5 72.4 73.9 74.0 72.9 73.8 86.8
ADP [41] 85.4 79.3 82.0 84.4 85.1 84.1 83.7 81.0 80.5 79.1 80.5
DiffPure [27] 72.5 69.8 71.8 70.1 72.2 73.4 73.5 72.7 70.2 69.4 89.3
CAD(ours) 95.8 93.4 83.8 94.7 95.0 93.7 93.7 95.2 95.4 93.7 95.8
$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 94.4 95.6 94.6 94.1 94.6 94.3 95.1 95.1 95.5 95.7 96.0
Table 1: Classification accuracy (%) against various attacks on CIFAR-10. Volumn ”Clean” represents the standard accuracy of clean images. $\mathrm{\textbf{CAD}}^{\dagger}$ denotes premium-CAD that is not constrained by the efficient memory and few-shot defense feedback.
Stage 0 1 2 3 4 5 6 7 8 9 -
Attack PGD VNIM FGSM BIM RFGSM MIM DIM NIM SNIM VMIM Clean
TRADES [44] 46.5 49.6 48.8 46.2 48.0 47.7 46.7 48.2 45.5 49.5 76.2
GAIRAT [45] 59.8 58.4 59.4 59.2 58.8 58.9 58.9 59.7 60.2 59.8 74.1
RPF [2] 56.4 51.3 55.8 56.6 54.4 52.6 53.9 53.3 49.9 50.1 69.4
DMAT [39] 64.3 66.2 66.4 65.9 65.4 64.4 64.1 65.8 65.1 65.6 75.9
ADP [41] 77.2 71.8 73.7 71.8 75.6 75.9 76.2 76.0 76.9 72.3 74.1
DiffPure [27] 53.2 52.4 53.7 55.8 53.9 53.4 52.6 54.7 53.0 52.3 73.6
CAD(ours) 83.7 82.0 76.2 82.5 83.5 82.0 81.9 82.2 83.4 81.9 88.1
$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 83.5 84.0 84.1 82.6 84.3 83.8 84.2 82.2 83.9 81.4 87.8
Table 2: Classification accuracy (%) on ImageNet-100.
@@ 5.1 Comparisons to Baselines
We compared CAD to other baseline methods in defending against 10 adversarial attacks. In our approach, the defense environment is dynamic, with attacks occurring at different stages. In contrast, the other baseline methods operate in a static defense environment that does not differentiate the order of attacks. We report the robust accuracy, which measures the performance on adversarial examples generated by attacks, and the standard accuracy, which measures the performance on clean data, separately for each method. We compare the defense performance of our method to other methods on CIFAR-10 and ImageNet-100. All the evaluation is conducted on the whole test set.
@@
As depicted in  Tab. 1 and Tab. 2, despite the modest scale of feedback, both CAD and premium-CAD demonstrate strong adaptability to each type of attack. When compared with the top baseline methods DMAT [39] and ADP [41], CAD outperforms them by approximately $10\%$ in CIFAR-10 and $6\%$ in ImageNet-100 for each attack. This underscores its capacity for continuous adaptation in response to emerging attacks. Another notable advantage of CAD is its ability to maintain high classification performance for clean images, achieving $95.8\%$ accuracy for CIFAR-10 and $88.1\%$ for ImageNet-100, while the testing accuracy of the target model stands at $96.4\%$ for CIFAR-10 and $89.8\%$ for ImageNet-100. This can be attributed to the generic representation learned by the weight estimator for adversarial examples, enabling it to effectively differentiate between clean and adversarial images. We observe that the accuracy against FGSM is relatively lower than that for other attacks, largely due to the distinct noise pattern of the one-step FGSM compared to other attacks.
To assess CAD’s performance in mitigating catastrophic forgetting, we present the average incremental accuracy in Fig. 3. The curve initially experiences a slight decline in the first three stages before stabilizing. This temporary dip is primarily attributable to FGSM’s influence, which pulls down the entire curve, and the adaptation to new attacks, resulting in a minor decrease in accuracy against previous attacks. Subsequently, the curve exhibits a modest increase due to the resemblance of other attacks to the initial attack PGD, thereby elevating the overall curve. Throughout the evaluation, the curves consistently hover around $92\%$ for CIFAR-10 and $80\%$ for ImageNet-10, underscoring CAD’s robust capacity to counter catastrophic forgetting.
@@
Note that premium-CAD outperforms all baseline methods. Given adequate defense feedback and memory resources, premium-CAD emerges as the superior choice.
@@ 5.2 Ablation Study
Figure 3: Average incremental accuracy (%) of CAD.
As argued, embedding reservation allows the defense model to be adapted more easily in the future and alleviates overfitting, prototype augmentation maintains the decision boundary of previous stages, and model ensemble ensures robust classification performance for both clean and adversarial images. Subsequently, we proceed to undertake an ablation study to validate the efficacy of each component.
As shown in Tab. 3 and Fig. 4 (Left), CAD without $L_{\mathrm{virtual}}$ and $L_{\mathrm{proto}}$ experiences a decrease of $13.2\%$ and $2.2\%$ respectivily in performance for adaptation. Furthermore, the absence of model ensemble results in the defense model’s inability to uphold classification accuracy for clean images alone. Additionally, we present curves illustrating the classification accuracy of each attack in Fig. 4 (Right). These curves demonstrate that the adaptation of CAD to FGSM is comparatively less effective than for other attacks, aligning with our earlier observations about FGSM in Sec. 5.1.
@@
Attack PGD VNIM FGSM BIM Clean
w/o $L_{\mathrm{virtual}}$ 95.6 86.6 70.6 78.5 95.8
w/o $L_{\mathrm{proto}}$ 95.7 93.7 81.6 88.2 95.8
w/o $f_{p}$ 95.8 93.4 83.8 94.7 55.9
CAD(ours) 95.8 93.4 83.8 94.7 95.8
Table 3: Classification accuracy (%) on CIFAR-10 for ablation.
Figure 4: Accuracy (%) in each stage on CIFAR-10. Left: average incremental accuracy of ablation. Right: classification accuracy against each attack.
@@ 5.3 Multiple Attacks in One Stage
In practice, it is common for multiple attacks to occur simultaneously within a single stage. To assess CAD’s adaptability in such scenarios, we evaluate the classification accuracy when faced with 1, 2, and 3 concurrent attacks. As presented in Tab. 4, when the number of attack increase to $2$ and $3$, the classification accuracy against these attacks still remain at a high level: $92.8\%$ and $92.5\%$ at stage $3$ respectively. The results presented in Tab. 4 indicate that CAD continues to perform effectively even when confronted with multiple simultaneous attacks.
@@
$\mathrm{number}=1$ 95.8 93.4 83.8 94.7
$\mathrm{number}=2$ 95.8 88.5 93.5 92.8
$\mathrm{number}=3$ 95.8 89.5 93.6 92.5
Table 4: Classification accuracy (%) on CIFAR-10 for different number of attacks in one stage.
@@ 5.4 Number of Shots for Adaptation
As indicated in Principle 2, the defense adaptation relies on few-shot defense feedback. To investigate the influence of the amount of feedback $K$ per class, we conducted tests with $K=15,10,5,1$. The results presented in Tab. 5 illustrate that defense performance diminishes as $K$ decreases. Nevertheless, even with reduced $K$, CAD continues to efficiently defend against adversarial attacks ($90.2\%$ when $K=1$ at stage $3$), highlighting its strong capability in few-shot scenarios.
@@
Attack PGD VNIM FGSM BIM
$K=15$ 95.8 93.7 81.2 94.8
$K=10$ 95.8 93.4 83.8 94.6
$K=5$ 95.8 93.2 80.8 93.7
$K=1$ 95.8 92.8 77.5 90.2
Table 5: Classification accuracy (%) on CIFAR-10 with different feedback scale $K$.
@@ 5.5 Memory Allowance for Adaptation
As outlined in Principle 3, we need to employ memory efficiently for continual defense. In order to demonstrate CAD’s effectiveness in conserving memory, we conducted tests to measure the occupied memory space and model size at each stage. As depicted in Tab. 6, CAD requires neither a large cache footprint (with only a 25KB growth per stage) nor an expansion of the model size. This underscores our approach’s ability to utilize efficient memory.
@@
Table 6: Comparison of the Cache Size and Model Size (Byte) between CAD and premium-CAD on CIFAR-10.

### 13 Mar 2024 v2

@@ 4.4 Model Ensemble for Clean and Adversarial Image Classification
Input: Benign dataset $\mathcal{D}_{\mathrm{train}}$, initial attack $A_{0}$, target model $f_{t}$, and few-shot feedback $\mathcal{A}_{\mathrm{train},K}^{i}$ generated by attack $A_{i}$ in stage $i=1,\cdots,T$.
@@ 5 Experiment
To validate the performance of CAD against various adversarial attack methods in the dynamic scenario, we conduct extensive empirical studies on two datasets and compare CAD to baselines from three research streams. For evaluation and analysis, we use two metrics in this section: (1) Classification accuracy against each attack after adaptation for defense performance evaluation. (2) Average accuracy of all attacks occurring within and before each stage, i.e. average incremental accuracy [27], for assessing the knowledge retention of the model.
@@
We choose PGD-$l_{\infty}$ [23] as the initial attack. After that, we select 8 adversarial attack methods under $l_{\infty}$: BIM [18], RFGSM [34], MIM [6], DIM [40], NIM [21], SNIM [21], VNIM [38] and VMIM [38] to compose the attack pool in defense performance evaluation. The perturbation magnitude of these attacks is set to $\epsilon=8/255$ for CIFAR-10 and $\epsilon=4/255$ for ImageNet-100 following [44]. As for evaluating the capability of CAD in migrating catastrophic forgetting, $l_{\infty}$ attack VMIM [38] and SNIM [21], $l_{2}$ attack CW [1] and DeepFool [25], $l_{1}$ attack EAD [2] and EADEN [2], and $l_{0}$ attack OnePixel [31] and SparseFool [24] compose the attack pool. Details of attack parameters are given in the supplementary.
Stage 0 1 2 3 4 5 6 7 8 -
Attack PGD SNIM BIM RFGSM MIM DIM NIM VNIM VMIM Clean
None-defense 0.6 1.3 0.0 0.0 0.0 0.0 0.2 0.0 0.0 96.4
TRADES [44] 55.8 60.2 57.4 55.9 55.8 57.1 57.1 58.9 57.3 84.8
JEM [10] 55.8 62.0 46.5 52.3 52.4 51.1 51.4 57.0 48.2 91.7
PAT [19] 45.5 45.1 43.8 45.5 45.3 45.2 45.2 43.9 45.4 69.3
GAIRAT [45] 66.9 70.4 66.4 66.9 66.7 66.5 66.5 68.6 66.4 89.4
FastAdv [14] 37.1 39.6 36.8 37.0 37.7 37.7 54.6 36.4 37.6 75.2
RPF [5] 60.9 52.8 52.2 58.9 54.9 57.6 57.5 53.5 52.6 83.5
DMAT [39] 71.7 76.0 73.1 71.7 71.7 73.0 73.0 75.5 73.0 92.4
EBM [12] 75.2 72.9 73.0 76.6 76.5 72.4 73.9 74.0 73.8 86.8
ADP [41] 85.4 80.5 79.3 84.4 85.1 84.1 83.7 81.0 79.1 80.5
DiffPure [26] 72.5 70.2 69.8 70.1 72.2 73.4 73.5 72.7 69.4 89.3
CAD(ours) 94.5 95.4 93.4 94.7 95.0 93.7 93.7 95.2 93.7 95.8
$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 94.4 95.5 95.6 94.1 94.6 94.3 95.1 95.1 95.7 96.0
Table 1: Classification accuracy(%) against various attacks($l_{\infty}$) on CIFAR-10. Volumn ”Clean” represents the standard accuracy of clean images. $\mathrm{\textbf{CAD}}^{\dagger}$ denotes premium-CAD that is not constrained by efficient memory and few-shot defense feedback.
@@
Both the target model and defense model use ResNet as the backbone network in our method. Specifically, we use WideResNet-28-10 [42] for CIFAR-10 following [10] and ResNet-50 [11] for ImageNet-100 following [19]. A 4-layer ConvNet is used as the weight estimator model. The cosine classifier [8] is adopted as the classifier of the defense model. We utilize Torchattacks [15] for generating adversarial images. The amount of feedback per class is set to $K=10$. The number of stages is set to $T=8$ since there are 9 attacks (PGD is the initial attack). Parameters of Beta distribution are set to $\alpha=\beta=2$ and the trade-off parameter for $\mathcal{L}_{0}$ is set to $\gamma=0.01$ following [47]. Epochs for training $f_{0}$ are set to $100$ and fine-tuning epochs are set to $4$.
Defense Baselines
We first compare to adversarial training methods TRADES [44], JEM [10], GAIRAT [45], FastAdv [14], RPF [5] and DMAT [39], and purification methods EBM [12], ADP [41] and DiffPure [26] in defense performance evaluation. As for evaluating the capability of CAD in migrating catastrophic forgetting, we compare to CL methods EWC [16], SSRE [51], BEEF [36], and FACT [47]. Besides, we propose another variant of our method denoted as premium-CAD in the case of no constraint of memory and no concerns about the amount of feedback. In premium-CAD, we train additional attack-specific defense models using abundant feedback in each stage, and ensemble them with the target model by the weight estimator. Details about the implementation of defense baseline and premium-CAD are given in the supplementary.
@@ 5.1 Comparisons to Baselines
To evaluate the defense performance, we first compared CAD to defense baseline methods in defending against 9 adversarial attacks. For CAD, the defense environment is dynamic, with attacks occurring at different stages. The other baseline methods operate in a static defense environment that does not differentiate the order of attacks. All the evaluation is conducted on the whole test set of CIFAR-10 and ImageNet-100. Classification accuracy is reported in Tab. 1 and Tab. 2.
Stage 0 1 2 3 4 5 6 7 8 -
Attack PGD SNIM BIM RFGSM MIM DIM NIM VNIM VMIM Clean
None-defense 0.2 0.0 0.0 0.0 0.0 0.0 0.1 0.0 0.0 89.8
TRADES [44] 46.5 45.5 49.6 46.2 48.0 47.7 46.7 48.2 49.5 76.2
GAIRAT [45] 59.8 60.2 58.4 59.2 58.8 58.9 58.9 59.7 59.8 74.1
RPF [5] 56.4 49.9 51.3 56.6 54.4 52.6 53.9 53.3 50.1 69.4
DMAT [39] 64.3 65.1 66.2 65.9 65.4 64.4 64.1 65.8 65.6 75.9
ADP [41] 77.2 76.9 71.8 71.8 75.6 75.9 76.2 76.0 72.3 74.1
DiffPure [26] 53.2 53.0 52.4 55.8 53.9 53.4 52.6 54.7 52.3 73.6
CAD(ours) 83.7 83.4 82.0 82.5 83.5 82.0 81.9 82.2 81.9 88.1
$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 83.5 83.9 84.0 82.6 84.3 83.8 84.2 82.2 81.4 87.8
Table 2: Classification accuracy(%) against various attacks($l_{\infty}$) on ImageNet-100.
Stage 0 1 2 3 4 5 6 7 8
Attack PGD SNIM VNIM CW DeepF. EAD EADEN OnePixel SparseF.
None-defense 0.6 1.3 0.0 0.0 8.6 0.0 0.2 39.1 29.7
EWC [16] 95.5 68.2 67.0 23.7 23.6 23.7 26.9 25.9 22.2
SSRE [51] 89.1 82.5 78.1 65.8 58.4 50.6 47.5 42.4 37.8
BEEF [36] 94.9 83.1 84.7 82.3 77.9 74.2 73.8 72.6 64.5
FACT [47] 94.5 87.2 86.4 81.5 80.9 78.6 78.4 78.2 74.6
CAD(ours) 94.5 89.1 90.5 84.1 84.8 81.1 82.1 82.8 78.2
$\mathrm{\textbf{CAD}}^{\dagger}$(ours) 94.4 95.5 95.1 95.2 94.9 95.3 94.6 96.0 95.2
Table 3: Continual defense performance (Average incremental accuracy(%)) against attacks under different norms $l_{p=0,1,2,\infty}$ on CIFAR-10. DeepF. and SparseF. denote DeepFool [25] and SparseFool [24] respectively.
@@
As depicted in Tab. 1 and Tab. 2, despite the small scale of feedback, both CAD and premium-CAD demonstrate strong adaptability to each type of attack. When compared with the top baseline methods DMAT [39] and ADP [41], CAD outperforms them by approximately $10\%$ on CIFAR-10 and $6\%$ on ImageNet-100 for each attack. This underscores its capacity for continuous adaptation in response to emerging attacks. Another notable advantage of CAD is its ability to maintain high classification performance for clean images, achieving $95.8\%$ accuracy on CIFAR-10 and $88.1\%$ on ImageNet-100, while the target model stands at $96.4\%$ on CIFAR-10 and $89.8\%$ on ImageNet-100. This can be attributed to the generic representation learned by the weight estimator for adversarial examples, enabling it to effectively differentiate between clean and adversarial images.
To assess the effectiveness of CAD in retaining knowledge, we compare it to CL baseline methods through experiments on CIFAR-10 with attacks under different norms $l_{p=0,1,2,\infty}$. All baseline methods are functioned within our dynamic scenario. As depicted in Tab. 3, both CAD and premium-CAD maintain robust defense performance despite changes in attack types. However, as the variety of attack types increases, the defense performance gradually diminishes. This indicates that defending against an expanding array of attack types presents a more challenging task for continual defense. Compared with the top baseline method FACT [47], CAD consistently outperforms it by an average of approximately $3.4\%$, highlighting CAD’s effectiveness in mitigating catastrophic forgetting.
@@
It is noteworthy that premium-CAD surpasses all baseline methods. Given sufficient defense feedback and memory resources, premium-CAD emerges as the superior choice.
(a) PGD
(b) CW
(c) EAD
(d) OnePixel
Figure 3: T-SNE diagrams of features extracted from adversarial test data encoded by the defense model after all stages. Each color represents a class of CIFAR10. After several steps of adaptation, CAD retains the capability to categorize the attacked data.
@@ 5.2 Ablation Study
As argued, embedding reservation facilitates easier adaptation of the defense model and mitigates overfitting. Prototype augmentation preserves the decision boundary of previous stages, while model ensemble ensures robust classification performance for both clean and adversarial images. Therefore, we proceed to conduct an ablation study using 4 attacks PGD($l_{\infty}$), CW($l_{2}$), EAD($l_{1}$), and
@@
Attack PGD CW EAD OnePixel Clean
w/o $L_{\mathrm{virtual}}$ 94.5 82.4 76.1 71.8 95.8
w/o $L_{\mathrm{proto}}$ 94.4 89.5 80.0 78.5 95.8
w/o $f_{p}$ 94.5 89.9 86.6 85.3 55.9
CAD(ours) 94.5 89.9 86.6 85.3 95.8
Table 4: Average incremental accuracy(%) on CIFAR-10 for ablation.
OnePixel($l_{0}$) to validate the efficacy of each component. As shown in Tab. 4, CAD without $L_{\mathrm{virtual}}$ and $L_{\mathrm{proto}}$ experiences a decrease of $13.5\%$ and $6.8\%$ respectivily in the last stage. The absence of model ensemble leads to the inability to uphold performance on clean images.
Additionally, we provide T-SNE diagrams of features extracted from adversarial test data in Fig. 3 to visualize the defense capability of CAD after adaptation. These features are encoded by the defense model after the last stage of adapting to OnePixel attack. From (a) to (d) in Fig. 3, the clustering of features becomes less and less discernible, indicating that the difficulty of few-shot adaptation increases with the introduction of new attack types. After adaptation, CAD maintains its defense performance against previous attacks, illustrating its ability to mitigate catastrophic forgetting as demanded in Principle 1.
@@ 5.3 Multiple Attacks in One Stage
In practice, it is common for multiple attacks to occur simultaneously within a
@@
$\mathrm{number}=1$ 94.5 95.4 93.4 94.7
$\mathrm{number}=2$ 94.5 94.4 93.5 92.8
$\mathrm{number}=3$ 94.5 93.9 93.6 92.5
Table 5: Classification accuracy(%) about the number of attacks($l_{\infty}$) in one stage.
single stage. To assess CAD’s adaptability in such scenarios, we evaluate the classification accuracy when faced with 1, 2, and 3 concurrent attacks on CIFAR-10. As presented in Tab. 5, when the number of attacks increases to $2$ and $3$, the classification accuracy against these attacks still remains at a high level: $92.8\%$ and $92.5\%$ at stage $3$ respectively. The results indicate that CAD continues to perform effectively even when confronted with multiple simultaneous attacks.
@@ 5.4 Number of Shots for Adaptation
As indicated in Principle 2, the defense adaptation relies on few-shot defense feedback. To investigate the influence of the amount of feedback $K$ per class, we
@@
Attack PGD CW EAD OnePixel
$K=15$ 94.5 90.0 88.4 85.5
$K=10$ 94.5 89.9 86.6 85.3
$K=5$ 94.5 89.4 86.1 84.6
$K=1$ 94.5 87.9 84.6 83.4
Table 6: Average incremental accuracy(%) with different feedback scale.
conducted tests with $K=15,10,5,1$ on CIFAR-10. The results presented in Tab. 6 illustrate that defense performance diminishes as $K$ decreases. Nevertheless, even with reduced $K$, CAD continues to efficiently defend against adversarial attacks ($93.4\%$ when $K=1$ at stage $3$), highlighting its strong capability in few-shot scenarios.
@@ 5.5 Memory Allowance for Adaptation
As outlined in Principle 3, we need to employ memory efficiently for continual defense. In order to demonstrate CAD’s effectiveness in conserving memory, we conducted tests to measure the occupied memory space and model size at each
@@
Table 7: Comparison of the Cache Size and Model Size (Byte) between CAD and premium-CAD.
stage on CIFAR-10. As depicted in Tab. 7, CAD requires neither a large cache footprint (with only a 25KB growth per stage) nor an expansion of the model size. This underscores our approach’s ability to utilize efficient memory.